OpenStudy (anonymous):

A student recorded the following buret readings during a titration of a base with an acid: Standard 0.100 M HCl: Initial Reading was 9.08 mL, Final Reading was 19.09 mL Unknown KOH: Initial Reading was 0.55 mL, Final Reading was 5.56 mL. Calculate the molarity of the KOH and show all work. Record the answer to the correct number of significant figures.

OpenStudy (anonymous):

First, write and balance the equation. Secondly, consider the molar ratios, which are the same as volume ratios. so for this case, you could situ that the amount of HCl reacted was 19.09-9.08=10.01mL and that of KOH was 5.56-0.55=5.01mL. Going onto the problem we get HCl + KOH - KCl + H2O we need to find moles of HCl under the formula n=CV/1000 or 0.1*10.01/1000= 0.0001001 mol, which are the same for KOH. Still in the same formula, we make C the subject for KOH and C=1000n/V= 1000*0.0001001/5.01=0.019M